OpenStudy (anonymous):

please help me solve this problem : -3ax^6(-5ax^7+9x^2)

OpenStudy (anonymous):

are you supposed to multiply ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\((-3ax^6)\times (-5ax^7)=15a^2x^{13}\) by the laws of exponents

OpenStudy (anonymous):

and \((-3ax^6)\times (9x^2)=-27ax^8\)

OpenStudy (anonymous):

so your answer is \[15a^2x^{13}-27ax^8\]

OpenStudy (anonymous):

thank you so much